### v2

# Enhancing Financial Time-Series Forecasting with Retrieval-Augmented Large Language Models

Abstract.
Stock movement prediction, a critical task in financial time-series forecasting, relies on identifying and retrieving key influencing factors from vast and complex datasets. However, traditional text-trained or numeric similarity-based retrieval methods often struggle to handle the intricacies of financial data. To address this, we propose the first retrieval-augmented generation (RAG) framework specifically designed for financial time-series forecasting. Our framework incorporates three key innovations: a fine-tuned 1B large language model (StockLLM) as its backbone, a novel candidate selection method enhanced by LLM feedback, and a training objective that maximizes the similarity between queries and historically significant sequences. These advancements enable our retriever, FinSeer, to uncover meaningful patterns while effectively minimizing noise in complex financial datasets. To support robust evaluation, we also construct new datasets that integrate financial indicators and historical stock prices. Experimental results demonstrate that our RAG framework outperforms both the baseline StockLLM and random retrieval methods, showcasing its effectiveness. FinSeer, as the retriever, achieves an 8% higher accuracy on the BIGDATA22 benchmark and retrieves more impactful sequences compared to existing retrieval methods. This work highlights the importance of tailored retrieval models in financial forecasting and provides a novel, scalable framework for future research in the field.
Retrieval-augmented generation, Financial Time-series forecasting, Stock Movement Prediction, Large Language Model
copyright: noneccs: Information systems Top-k retrieval in databasesccs: Information systems Similarity measuresccs: Information systems Novelty in information retrieval
1. Introduction
Financial time-series forecasting plays a pivotal role in ensuring market stability and efficiency, directly impacting critical areas such as investment strategies, risk management, and the formulation of economic policies (Fama and French, 2000). One of the most prominent tasks in this domain is stock movement prediction (Xie et al., 2023a, 2024b, 2024a), which focuses on forecasting the direction of price changes, i.e. whether a stock will rise or fall. Accurate predictions of financial metrics, including stock movements, interest rates, and economic indicators, are essential for a wide range of stakeholders, from individual investors to large financial institutions. However, the inherent complexity and volatility of financial markets pose significant challenges to forecasting, necessitating the use of advanced methodologies to effectively analyze vast and often noisy datasets.
Traditional stock movement prediction methods (Feng et al., 2018; Qin et al., 2017) relied on analyzing sequences of stock prices and financial indicators [^1] Financial indicators (Jansen, 2020) are professional market analysis metrics derived from price data, where their values or crossover points signal fluctuations in the stock market, indicating rising or falling trends. However, relying on one or two indicators can introduce bias, and if an indicator only provides meaningful information at crossover points, it may fail to offer valuable insights during periods without crossings. to identify patterns, but they often failed to capture the semantic connections between these sequences, a task at which large language models (LLMs) excel (Huang and Chang, 2022). However, LLM-based methods (Xie et al., 2023b, 2024a, a) have primarily shifted focus to leveraging textual data such as news and tweets, often using only the past several days’ closing prices as reference while overlooking the rich insights embedded in historical time-series data (Bustos and Pomares-Quimbaya, 2020). This highlights the need to effectively integrate financial time-series data with LLMs. Directly working with such data presents a challenge due to its vast scale, encompassing both the breadth of influencing variables and the depth of historical trends. To address this, a retrieval-based approach is necessary to efficiently sift through extensive time-series data, extracting meaningful information to enhance LLMs’ ability to deliver more accurate and robust stock movement predictions.
Although retrieval-augmented generation (RAG) methods have been applied for various tasks (Joshi et al., 2024; He et al., 2024; Cui et al., 2024), it is challenging to directly apply them to financial time-series retrieval. First, most embedding-based retrieval methods (retrievers) (Su et al., 2023; Li and Li, 2023; Xiao et al., 2023; Zhang et al., 2023; Wang et al., 2023) are trained on textual data and struggle to process numeric time-series data, which lacks explicit semantic information. Second, distance-based retrieval methods like Dynamic Time Warping (DTW) (Yang et al., 2024) focus on numeric similarities and fail to capture deeper semantic relationships or contextual nuances essential for stock movement prediction. These limitations underscore the need for a more tailored retrieval framework designed for financial time-series data.
[…]
The framework of FinSeer, illustrating its components and workflow.
Figure 1. Overview of our time-series RAG framework.
To address these challenges, we introduce a retrieval-augmented framework for stock movement prediction, featuring the first dedicated retriever for financial time-series forecasting, Financial Time-Series Retriever (FinSeer). Unlike previous RAG methods, which rely on pre-trained encoders or distance-based metrics, our framework is the first to train a dedicated retriever tailored for continuous and complex temporal sequences. The whole retrieval framework is shown in Figure 1.
To address the limitations of existing datasets, which often contain only price data, and to better align with professional financial analysis, we construct new datasets by integrating 20 groups of financial indicators. These indicators provide critical insights into market behavior that stock prices alone cannot capture [^2] For instance, the overbought area in the KDJ indicator group (Wu and Diao, 2015), where $K>80$, $D>70$, and $J>90$, signals a potential falling trend.. We begin by selecting high-trade-volume U.S. stocks, and the indicators are chosen through domain expert consultation and mutual information analysis. These indicators are then serialized and divided into train, valid and test sets. By training the retriever on datasets enriched with diverse financial indicators, our retrieval framework incorporates financial domain expertise, enabling it to interpret the implications of trending signals that uncover deeper patterns beyond surface-level trend similarities. The testing set, designed to mimic real-world economic analysis, ensures a robust evaluation of model performance in practical financial scenarios.
To train our retriever to identify semantically related sequences and align the retriever’s priorities with those of the LLM, we introduce a novel retrieval mechanism with three key innovations: a new backbone LLM, a novel method for selecting positive and negative candidates, and a new training objective. First, to activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA3.2-1B Instruct) (Grattafiori et al., 2024), creating StockLLM as the backbone model of our retriever. By using a smaller LLM, we establish a more challenging experimental setup, ensuring performance improvements are attributable to FinSeer rather than the LLM’s capacity. Next, StockLLM identifies time-series segments that enhance the generative process, feedbacks from StockLLM are used to select the most beneficial sequences that could lead the model to make the correct decision as positive candidates and the least beneficial as negative candidates. This could bridge the gap between the retriever and the LLM and ensure that retrieved data aligns with the LLM’s forecasting priorities. Finally, the training objective, inspired by Zhang et al. (2023), ensures the retriever distinguishing historically significant sequences (positive candidates) from noisy sequences (negative candidates) by maximizing the similarity between the query and candidate sequences. This prioritizes meaningful patterns while minimizing irrelevant data, outperforming traditional methods that struggle to separate signal from noise in complex financial time series.
Based on FinSeer, the inference process of our RAG framework consists of three steps: (1) using FinSeer to retrieve relevant sequences from the candidate pool, (2) incorporating these sequences into StockLLM’s input context, and (3) combining this enriched context with the original query to perform forecasting. This approach effectively integrates salient historical patterns with current data, improving prediction accuracy.
To thoroughly evaluate the performance of our RAG framework and the retriever, we rely solely on temporal information, excluding additional textual data like news. For the RAG framework evaluation, StockLLM serves as the backbone model, integrated with different retrieval methods for comparison. Experimental results demonstrate that using our RAG framework outperforms bare StockLLM, indicating that the RAG framework enhances LLM performance by incorporating relevant time-series data as financial context. However, RAG with retrieval models trained on text data does not always improve performance compared to bare StockLLM. In some cases, performance even declines. This suggests that these retrieval models cannot effectively identify beneficial time-series data that could enhance the model’s predictions, which highlights a fundamental gap between text-based retrieval and time-series data. Furthermore, RAG framework with FinSeer consistently surpasses all other retrieval models, such as Instructor (Su et al., 2023), BGE (BAAI General Embedding) (Xiao et al., 2023), LLM-Embedder (Zhang et al., 2023), UAE (Li and Li, 2023), and E5-mistral-7b-instruct (Wang et al., 2023), demonstrating the effectiveness of our trained FinSeer in identifying relevant time-series financial data to enhance stock movement prediction. Further analysis of retrieval results reveals that FinSeer uniquely identifies relevant financial indicators that enhance LLM performance, while other retrievers predominantly select sequences based on superficial similarities, such as close price and adjusted close price. This behavior underscores a critical limitation of text-trained models: they focus on surface-level patterns and fail to discern truly valuable time-series information. In contrast, FinSeer effectively bridges this gap, demonstrating its ability to uncover underlying relevance in financial data.
Our contributions are summarized as follows:
  1. (1)
    We propose the first RAG framework for financial time-series forecasting, in which the retrieval mechanism features three key innovations: a fine-tuned 1B parameter LLM (StockLLM) as the backbone, a novel candidate selection method leveraging LLM feedback to identify beneficial sequences, and a training objective that maximizes similarity between queries and historically significant sequences, enabling the retriever to uncover meaningful patterns while minimizing noise in complex financial data.
  2. (2)
    We construct new training and testing datasets to address the limitations of existing ones containing only price data, and better align with real-world economic analysis. Besides price data, these datasets integrate financial indicators that signal upward or downward trends, carefully selected through domain expert consultation and mutual information analysis. By using new training datasets, we train the retriever to uncover deeper patterns beyond surface-level trends during training. While the testing set, designed to mimic real-world scenarios, ensures robust evaluation of model performance in practical financial applications.
  3. (3)
    We propose a new retriever, FinSeer, specifically designed for time-series forecasting, trained on newly constructed datasets that integrate financial indicators alongside historical stock prices. By leveraging LLM feedback and prioritizing historically significant sequences, FinSeer uncovers deeper patterns beyond surface-level trends, setting it apart from traditional methods.
  4. (4)
    Experimental results demonstrate that our RAG framework outperforms bare StockLLM and random retrieval, highlighting its effectiveness, while our retriever, FinSeer, surpasses existing RAG methods in stock movement prediction across three datasets, achieving an 8% higher accuracy on BIGDATA22 compared to a general-purpose LLM-feedback-based retriever. Furthermore, FinSeer retrieves more dynamic and impactful sequences, showcasing its ability to identify and leverage the most relevant patterns for forecasting.
2. Problem Definition and Goals
Retrieval-augmented financial time-series forecasting (Sezer et al., 2020) involves predicting future values or trends ($G$) based on a given query sequence ($q$) and a set of retrieved historical sequences ($c$). These sequences are collected over time at regular intervals. In the retrieval process, the goal of the retrieval model ($R$) is to efficiently identify and extract the most useful historical sequences from a large pool of candidates. By providing relevant context, the retrieval model enhances the forecasting model’s ability to make accurate and reliable predictions.
Stock Movement Prediction Task. In the specific task of stock movement prediction (Xu and Cohen, 2018), the problem is framed as a binary classification task: predicting whether a stock’s price will rise or fall on the next trading day. Given a query sequence $q$, which represents the stock’s price over the previous $t$ days, the model retrieves relevant sequences as context and predicts the stock’s movement $M_{q,d}$ for the next trading day $d$. Table 1 defines the major symbols used in this paper.
Table 1. The definition of symbols.
General Time-series Symbol Stock Movement Prediction Symbol Definitions
$q$ $q=\{q_{d-t},...,q_{d-1}\}$ The query time-series data. In stock movement prediction, $q$ refers to the query stock price sequence of length $t$, containing stock price data from trading day $d-t$ to trading day $d-1$.
$G(q)$ $G(q,d)\in\{\textit{rise},\textit{fall}\}$ The final output $G$ given the query $q$. In stock movement prediction, $G(q,d)$ shows the generation $G$ of the query stock $q$ on the query trading day $d$, belonging to rise or fall.
$P(c)=LLM(O|q,c)$ $P(c)=LLM(M_{d}|q,c)$ The probability $P$ of the LLM to generate an accurate output $O$ given the query sequence $q$ and a candidate sequence $c$. In stock movement prediction, $P(c)$ refers to generating the accurate movement $M$ on the query trading day $d$.
=$\mathbb{C_{P}}=\{c_{i}\mid i=1,\ldots,k\}$ The set of top-k retrieved sequences as positive examples, where $P(c_{i})\geq P(c_{i+1})$.
=$\mathbb{C_{N}}=\{c_{i}\mid i=k+1,\ldots,n\}$ The set of negative retrieved sequences, where $P(c_{i})\geq P(c_{i+1})$.
$w_{i}$ The soft weight of the $i^{th}$ retrieved sequences, where $w_{i}=P(c_{i}),i=1,...,k$.
$R$ The retrieve model.
Rise/Fall Threshold Settings. To classify daily movements as rise or fall, we first calculate returns $R_{t}$, which represents the percentage change in the closing price over consecutive days.
[…]
Following Yoo et al. (2021) and  Soun et al. (2022), we classify the movement as rise if return exceeds 0.55, fall if it is below -0.5. In line with previous experimental settings (Xie et al., 2023b, 2024a), we do not evaluate freeze cases. However, we include sequences with $R_{t}\in[-0.5,0.55]$ as freeze candidates in the candidate pool, ensuring a diverse and comprehensive set of historical data for context.
$M_{q,d}&=\begin{cases}\text{rise},&R_{t}>0.55\\ \text{fall},&R_{t}<-0.5\\ \text{freeze},&-0.5\leq R_{t}\leq 0.55\end{cases}\end{aligned}$
Rationale for Unbalanced Thresholds. The asymmetrical thresholds for classifying daily movements (0.55 for rises and -0.5 for falls) reflect the inherent dynamics of stock market behavior. Stock prices typically rise gradually due to sustained investor optimism but fall sharply during panic selling or profit-taking. The stricter rise threshold prevents minor upward fluctuations from being misclassified as significant increases, while the more lenient fall threshold ensures meaningful downward trends are captured. This approach aligns with market realities, improving the reliability of movement classifications.
3. The RAG Framework
To identify valuable sequences from vast amounts of stock data and conduct stock movement prediction, we propose a novel RAG framework with new datasets and a novel retrieval mechanism with a new backbone LLM. First, we construct three datasets that integrate financial indicators alongside historical stock prices, addressing the limitations of existing datasets and better aligning with real-world economic analysis. Second, we introduce StockLLM, a fine-tuned 1B parameter LLM, as the backbone to ensure instruction-following capabilities and establish a challenging experimental setup. Finally, we design a novel retrieval mechanism that leverages LLM feedback to select positive and negative candidates and employs a new training objective to prioritize historically significant sequences. This comprehensive framework bridges the gap between retrieval models and LLMs, enhancing the time-series forecasting process.
3.1. Dataset Construction
To address the limitations of existing datasets and better align with real-world economic analysis, we construct new datasets by integrating financial indicators alongside historical stock prices, enabling deeper insights into market behavior that stock prices alone cannot capture. This section details stock selection, raw data collection (Collection of Basic Price Indicators) and financial indicator selection process(Calculation and Selection of Financial Indicators). Specifically, our datasets are built using high-trade-volume U.S. stocks, while the indicators are selected through domain expert consultation and mutual information analysis.
Stock Selection. To construct our datasets, we select high-trade-volume U.S. stocks across three periods: 2014-2015, 2017-2018, and 2022-2023. The first two periods align with two benchmark datasets: for 2014-2015, we use the same stocks as the ACL18 dataset (Xu and Cohen, 2018), and for 2017-2018, we use the same stocks as the BIGDATA22 dataset (Soun et al., 2022). To incorporate more recent data, we manually select high-trade-volume stocks from 2022-2023 to create the STOCK23 dataset. This ensures our training and evaluation reflect recent market conditions and provide a robust benchmark for stock movement prediction.
Collection of Basic Price Indicators. Stock data is collected using the Yahoo Finance API (Xu and Berkely, 2014), which provides basic price indicators such as the opening price, highest price, lowest price, adjusted closing price, and trading volume for each trading day. The daily price movement is then calculated as the reference answer for model training and evaluation.
We then divide the stock data into training, validation, and test sets. To ensure temporal integrity and prevent information leakage, we partition the dataset by stocks rather than by time. This approach maintains consistent time ranges across splits while ensuring that stocks in the validation and test sets are entirely unseen during training. Unlike time-based splits, stock-based partitioning provides a more robust evaluation of the model’s ability to generalize to new stocks, better reflecting real-world scenarios where models are applied to unseen entities. Table 2 summarizes the stock counts for each split in our datasets.
Table 2. Dataset statistics.
Dataset Stock amount Trading dates
train valid test all sequences query sequences
ACL18 (Xu and Cohen, 2018) 33 5 33 2014.06.02-2015.12.31 2015.06.03-2015.12.31
BIGDATA22 (Soun et al., 2022) 22 3 22 2019.04.01-2020.12.31 2020.04.09-2020.12.31
STOCK23 24 3 24 2022.01.03-2023.12.31 2023.01.03-2023.12.31
Calculation and Selection of Financial Indicators. In our datasets, we integrate financial indicators to provide critical insights into market behavior that stock prices alone cannot capture. Following analyzing methods in the book Machine learning for trading (Jansen, 2020), we conduct an analysis using query sequences in all training sets to explore the relationship between commonly used financial indicators and returns. To measure the relationship between these indicators and returns, we utilize Mutual Information (MI), a non-linear measure of dependency that quantifies the amount of information one variable contains about another. Specifically, the MI between an indicator $X$ and returns $R_{t}$ is calculated as:
[…]
where $p(x,r)$ represents joint probability density function of $X$ (indicator) and $R_{t}$ (returns), $p(x)$ shows marginal probability density function of $X$, and $p(r)$ is marginal probability density function of $R_{t}$.
We compute MI scores for each indicator and select the top-20 indicator groups with the highest scores as our candidates, with the detailed score ranking provided in Appendix B.1. Additionally, an example of a data piece from our datastore is illustrated in Appendix A.1.
3.2. Sequence Serialization
Since stock movement prediction depends on the changes in related features rather than their exact values, we serialize stock prices and financial indicators into a time-series format. We use JSON to represent these sequences, as it has been demonstrated to effectively support LLMs in interpreting time-series data (Fang et al., 2024; Singha et al., 2023; Yin et al., 2023).
Query Sequence Serialization We select query sequences where the query date is at least one year after the start date of the corresponding split in the dataset. This ensures that each query has enough candidate sequences for retrieval. For example, in the ACL18 dataset, which spans from 2014-06-02 to 2015-12-31, the query sequences are defined as those from 2015-06-03 to 2015-12-31.
Each query sequence contains the adjusted close price from the previous five trading days along with basic stock information. The indicator is represented as a five-day sequence, with a one-day sliding window applied across trading days. For example, when inquerying about stock MO on 2015-06-02, the query sequences contains stock name, query date, last five trading dates and their corresponding adjusted close prices. The serialized sequence is shown below:
{ "query_stock": "MO",
"query_date": "2015-06-02",
"recent_date_list": ["2015-05-26", "2015-05-27", "2015-05-28", "2015-05-29", "2015-06-01"],
"adjusted_close_list": [29.669, 29.9872, 29.8657, 29.6227, 29.6227]}
Candidate Sequence Serialization As new information becomes available, the candidate pool dynamically incorporates sequences from the most recent trading days. For instance, when the query date is 2015-06-03, the candidate pool includes sequences from 2014-06-02 to 2015-06-02. When the query date advances to 2015-06-04, the sequence for 2015-06-03 is added to the candidate pool without the need for additional training. This demonstrates the scalability of our method.
For each candidate stock on a specific date, we select 6 basic price indicators and 20 groups of financial indicators. Each candidate sequence includes a typical financial indicator along with basic stock information. As the same as query serialization, the indicator is represented as a five-day sequence, with a one-day sliding window applied across trading days. For example, a candidate represent stock MO on date 2014-07-02, with highest price as its indicator. The sequence includes basic information containing stock name, date, movement, last five trading dates and their corresponding highest prices. The serialized sequence is shown below:
{ "candidate_stock": "MO",
"candidate_date": "2014-07-02",
"candidate_movement": "freeze",
"recent_date_list": ["2014-06-25", "2014-06-26", "2014-06-27", "2014-06-30", "2014-07-01"],
"high_list": [42.2, 42.0, 41.86, 42.28, 42.0]}
3.3. Retriever Training
To identify valuable information from vast amounts of sequences and uncover semantic relationships between them, we introduce a novel retrieval mechanism featuring three key innovations: a fine-tuned 1B parameter LLM (StockLLM) as the backbone, a candidate selection method leveraging LLM feedback to identify beneficial sequences and a training objective that maximizes similarity between queries and historically significant sequences, enabling the trained retriever to uncover meaningful patterns while minimizing noise in complex financial data. While inspired by (Zhang et al., 2023), our approach is uniquely tailored to financial time-series data, addressing the challenges of relevance assessment and alignment in a domain where traditional retrieval methods often fall short. This section details the training process of our framework, including the training of the LLM backbone, candidate scoring and selection, and the training of the retriever, which contains the training objective and knowledge distillation.
The LLM Backbone. To activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA 3.2-1B-Instruct) (Grattafiori et al., 2024) using the LoRA technique for efficient low-rank adaptation, resulting in StockLLM. By intentionally using a smaller backbone model, we establish a more challenging experimental setup, ensuring that performance improvements are attributable to FinSeer’s retrieval capabilities rather than the LLM’s capacity. The fine-tuning process is implemented using the LlamaFactory framework (Zheng et al., 2024), with the following configuration: a learning rate of 5e-5, a cosine scheduler, gradient accumulation over 8 steps, mixed-precision (fp16) training, and 5 epochs of training with regular evaluation to log metrics and select the best-performing checkpoint. This setup demonstrates the robustness of our framework in extracting meaningful insights even under constrained model size, highlighting the effectiveness of FinSeer in enhancing financial forecasting tasks.
Candidate Scoring. To determine whether a candidate sequence assists in predicting the movement of the query, we use LLM feedback to score each candidate. Specifically, for a given query $q$, we integrate the query sequence and each candidate sequence $c_{i}$ from the candidate pool as concurrent inputs to the LLM. The LLM outputs logits, which are unnormalized scores representing the model’s confidence for each possible class (e.g., ”rise” or ”fall”). These logits are transformed into probabilities $P(c)$ using the softmax function:
(4) $P(c)=\frac{e^{z_{c}}}{\sum_{j}e^{z_{j}}},$
where $z_{c}$ is the logit for the correct class (e.g., ”rise” if the true movement is upward) and $z_{j}$ represents the logits for all possible classes. The resulting probability $P(c)$ serves as the score for the candidate $c_{i}$ with respect to the query $q$.
Candidate Selection. We rank the candidate sequences in descending order based on their scores $P(c)$. The top-1 sequence is selected as a positive candidate, while the bottom 15 sequences are chosen as negative candidates. The sets of selected positive and negative sequences are denoted as $\mathbb{C_{P}}$ and $\mathbb{C_{N}}$, respectively.
Training Objective. Our retriever $R(q)$ is designed to intelligently distinguish between historically significant sequences $\mathbb{C_{P}}$ and noisy sequences $\mathbb{C_{N}}$. The training objective is to ensure that $R(q)$ prioritizes sequences from $\mathbb{C_{P}}$ while minimizing attention to those from $\mathbb{C_{N}}$. This is achieved by maximizing a similarity measure $sup(q,s)$ between the query sequence $q$ and candidate sequences $s$. Mathematically, the retriever’s objective is formulated as:
(5) $R(q)={\arg\max}_{s\in\mathbb{C_{P}}\cup\mathbb{C_{N}}}sup(q,s).$
By focusing on sequences that maximize $sup(q,s)$, the retriever ensures that the most informative and contextually relevant historical sequences are identified.
Knowledge Distillation. To leverage the scoring derived from the LLM, we employ knowledge distillation, which transfers knowledge from the teacher model (LLM) to the student model (retriever) by mimicking the teacher’s output distribution. This approach effectively captures nuanced patterns and predictions from the LLM. Specifically, we minimize the Kullback-Leibler (KL) divergence between the candidate distributions computed using the LLM’s rewards and those predicted by the embedding model. For each query $q$ and its candidate list {$\{\mathbb{C_{P}},\mathbb{C_{N}}\}$, we derive the LLM’s rewards for the candidates, denoted as $\{P(c_{i}),i=1,...,n\}$. To make these rewards suitable for distillation, we normalize them using a softmax function with temperature $\alpha$:
(6) $w_{i}=\text{softmax}_{R}\left(\frac{P(c_{i})}{\alpha}\right).$
The KL divergence is then computed as follows:
(7) $\min\sum_{c}-w_{i}\times\log\left(\frac{\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{c_{i}}\right\rangle/\tau\right)}{\sum_{{c}^{\prime}\in\mathbb{C}}\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{{c}^{\prime}}\right\rangle/\tau\right)}\right),$
where $\boldsymbol{e}_{q}$ and $\boldsymbol{e}_{c_{i}}$ are the embeddings of the query $q$ and candidate $c_{i}$, respectively, and $\tau$ is a temperature parameter. This loss function optimizes the similarity between the query embedding and the embeddings of the top-ranked candidates, enhancing the retriever’s ability to accurately predict stock price movements.
3.4. Inference
The inference process of our RAG framework, built on FinSeer, consists of three key steps to enhance stock movement prediction. First, FinSeer retrieves relevant sequences from the candidate pool, leveraging its ability to identify historically significant patterns. Second, these retrieved sequences are incorporated into StockLLM’s input context, enriching the model’s understanding of temporal dynamics. Finally, the enriched context is combined with the original query to generate forecasts. By integrating salient historical patterns with current data, this approach bridges the gap between past trends and present conditions, significantly improving prediction accuracy and robustness in real-world financial scenarios.
4. Experiment
4.1. Experimental Settings
Datasets. We evaluate the effectiveness of our RAG framework on the test sets of the three datasets described in Table 2, with ACL18 containing 2,876 query sequences, BIGDATA22 containing 2,868 queries, and STOCK23 containing 4,128 queries. These thousand-scale queries help mitigate random bias, ensuring a robust and reliable evaluation of model performance. To ensure a comprehensive evaluation, for each query sequence, we include all sequences containing financial indicators across all stocks in the test set (not limited to the same stock), with data available up to the query date, as potential candidates. No additional restrictions are imposed, enabling a robust assessment of the models’ performance in real-world financial forecasting scenarios.
Baselines. To evaluate the efficiency of our RAG framework, we choose bare StockLLM and random retrieval as our baselines. To evaluate our retriever FinSeer, we select five competitive RAG models from the top of the MTEB English Leaderboard as baselines: (1) Instructor-large (Su et al., 2023): A 335M instruction-finetuned text embedder that encodes sequences into 768-dimensional tensors. (2) UAE-large-v1 (Li and Li, 2023): A 335M ANGLE-optimized text embedding model that encodes sequences into 1024-dimensional tensors. (3) BGE-large-en-v1.5 (Xiao et al., 2023): A 335M general embedder pre-trained with RetroMAE (Xiao et al., 2022), encoding sequences into 1024-dimensional tensors. (4) LLM-Embedder (Zhang et al., 2023): A 109M embedder fine-tuned from BGE-large-en-v1.5, designed as a unified embedding model to support diverse retrieval augmentation needs for LLMs. It encodes sequences into 768-dimensional tensors. (5) E5-mistral-7b-instruct (Wang et al., 2023): A 7111M embedder initialized from Mistral-7B-v0.1 (Jiang et al., 2023) and fine-tuned on multilingual datasets, encoding sequences into 4096-di-mensional tensors.
Evaluation Metrics. We employ Accuracy (ACC) and Matthews Correlation Coefficient (MCC) (Matthews, 1975) to assess the performance of FinSeer and the baseline models on the stock movement prediction task. These metrics evaluate the performance of stock movement prediction based on the distribution of positive and negative samples.
4.2. Main Results
Table 3 summarizes the performance of stock movement prediction across the three datasets.
Table 3. Results of stock movement predictions using LLMs and retrieval models. The asterisk (*) indicates the LLM employed while using retrieval models.
Retrieving Methods ACL18 BIGDATA22 STOCK23
(+ StockLLM) ACC MCC ACC MCC ACC MCC
w/o Retrieval 0.496 -0.007 0.497 -0.008 0.509 0.021
Random Retrieve 0.481 -0.036 0.521 0.045 0.496 -0.004
DTW 0.518 0.045 0.505 0.021 0.492 -0.007
Instructor 0.496 -0.008 0.497 -0.006 0.505 0.010
UAE 0.482 -0.036 0.495 -0.008 0.494 -0.009
E5 0.510 0.028 0.503 0.008 0.499 -0.001
BGE 0.489 -0.019 0.487 -0.026 0.488 -0.014
LLM Embedder 0.499 0.000 0.459 -0.083 0.503 0.007
FinSeer 0.521 0.042 0.538 0.079 0.542 0.085
We analyze our framework and retriever separately to evaluate their contributions. First, our RAG framework outperforms both bare StockLLM and random retrieval across all datasets, demonstrating its effectiveness in enhancing stock movement prediction. Second, among all evaluated retrieval methods, FinSeer consistently achieves the best performance when integrated with StockLLM, showcasing its ability to retrieve valuable sequences that improve forecasting accuracy. Notably, FinSeer is the only retriever that achieves stable improvements over bare StockLLM on all datasets, while other methods exhibit inconsistent results, sometimes enhancing predictions and other times degrading them. Furthermore, our experiments reveal that incorporating a retrieval process does not always guarantee performance improvements, as retrieval errors can mislead the LLM and harm predictions. This underscores the critical need for retrieval models that align closely with the LLM’s objectives, a challenge that FinSeer effectively addresses through its tailored design.
Figure 2. Indicator occurrences of different RAG models on ACL18 dataset.
The minimal performance difference between the instruction-finetuned retriever (Instructor) and the no-retrieval baselines highlights the inherent challenge of time-series retrieval. Unlike text retrieval, time-series retrieval does not solely rely on task understanding, as candidate sequences often appear similar, making it difficult to distinguish valuable ones. The angle-optimized retriever (UAE) shows inconsistent performance across datasets, as its ability to differentiate similarity between query and candidate sequences does not fully align with the LLM’s perception of importance. Similarly, LLM Embedder, our backbone model trained with LLM feedback, struggles to generalize to time-series retrieval tasks, further underscoring the complexity of the problem.
The best-performing baseline is E5, which achieves competitive results despite its significantly larger model size and embedding dimensions (five times larger than FinSeer). However, FinSeer outperforms E5 on all datasets. This demonstrates that FinSeer successfully learns the preferences of LLMs and effectively enhances their time-series forecasting ability through retrieval, even under challenging conditions.
In summary, while RAG methods do not guarantee performance improvements due to the risk of retrieval errors, FinSeer stands out by consistently retrieving high-quality sequences that align with the LLM’s forecasting objectives, leading to superior results compared to other retrievers.
4.3. Ablation Study
Indicator Occurrences In this part, we analyze what indicators all RAG models retrieve. We calculate indicator occurrences on ACL18 dataset and the results are shown in Figure 2.
As shown in the figure, FinSeer is the only model that successfully extracts a diverse and comprehensive set of indicators while achieving superior performance. This clearly demonstrates its advanced temporal retrieval capabilities. Specifically, while other models like LLM Embedder and Instructor predominantly focus on basic indicators such as close price and adjusted close price, FinSeer effectively identifies and retrieves a wide range of technical indicators, including kdj crossover, MACD Histogram, Bollinger Bands, and various alpha factors. This richer set of retrieved indicators provides FinSeer with more comprehensive auxiliary information, enabling more accurate and reliable predictions.
Candidate Movement Correlation In this section, we investigate whether the LLM relies on the movement trends of retrieved sequences to make predictions. To this end, we compute the correlation between the movements of retrieved sequences and the LLM’s generated results. The detailed calculation process is provided in Appendix B.2.
The results in Table 4 reveal no significant correlation between the movement direction (rise or fall) of the retrieved candidate sequences and the final predictions of StockLLM. This finding indicates that StockLLM does not simply mirror the movement trends of the retrieved sequences but instead analyzes their specific content to infer the query’s movement direction. This ability highlights the LLM’s capacity to extract meaningful insights from complex time-series data.
Table 4. Correlation between the movements of retrieved sequences and LLM’s generated results.
Dataset RAG Model Correlation P-Value
ACL18 Instructor -0.098 0.000
UAE -0.068 0.000
E5 -0.108 0.000
BGE 0.175 0.000
LLM Embedder 0.014 0.457
FinSeer 0.054 0.004
BigData22 Instructor -0.072 0.000
UAE 0.009 0.633
E5 -0.040 0.032
BGE 0.093 0.000
LLM Embedder 0.060 0.001
FinSeer 0.077 0.000
STOCK23 Instructor -0.114 0.000
UAE -0.093 0.000
E5 -0.064 0.000
BGE 0.131 0.000
LLM Embedder 0.015 0.350
FinSeer 0.032 0.041
This observation underscores the critical role of retrieval quality in RAG models, as the relevance and informativeness of retrieved sequences directly influence prediction performance. Notably, StockLLM is a relatively small 1B parameter LLM, making this a particularly challenging setting. Despite this, our time-series-tailored RAG model, FinSeer, successfully enhances StockLLM’s performance by retrieving rich and relevant time-sensitive information. This demonstrates the effectiveness of our approach in improving prediction accuracy and showcases its potential for financial time-series forecasting.
4.4. Case Study
This case study illustrates the critical importance of alignment between the retriever and the LLM’s forecasting preferences in financial time-series analysis. We examine the stock XOM on 2015-06-25 from the ACL18 dataset, where the adjusted close price exhibited a pronounced downward trend. The query sequence is as follows:
{ "query_stock": "XOM",
"query_date": "2015-06-25",
"recent_date_list": ["2015-06-18", "2015-06-19", "2015-06-22", "2015-06-23", "2015-06-24"],
"adjusted_close_list": [58.0813, 57.8979, 57.8707, 57.8027, 57.5377]}
While multiple retrievers were evaluated, only FinSeer successfully enabled StockLLM to predict the correct movement as a fall. Specifically, FinSeer retrieved five diverse indicators—close price, adjusted close price, alpha021, alpha054, and the highest price—providing a comprehensive view of the stock’s behavior. Alpha021 identifies trends based on short- and long-term price averages and volume conditions, while alpha054 combines price and volume rankings to assess performance within a specific time window. These indicators allowed StockLLM to accurately assess whether the downward trend would persist or reverse, demonstrating the value of retrieving contextually relevant and diverse features.
In contrast, other retrievers, such as Instructor, BGE, LLM Embedder, and E5, extracted sequences dominated by close or adjusted close prices, all reflecting similar downward trends. While these sequences aligned with the current trend, they failed to provide actionable insights for forecasting future movements, leading StockLLM to misinterpret them as noise and incorrectly predict a rise. Similarly, UAE retrieved sequences indicating overbought and oversold conditions, including three rise and two freeze trends. Although overbought signals often suggest a potential downturn, the retrieved sequences themselves exhibited rising or frozen trends, confusing StockLLM and resulting in an erroneous prediction. This case study underscores the superiority of FinSeer in retrieving meaningful and diverse indicators that align with the LLM’s forecasting logic, enabling more accurate and reliable predictions.
4.5. Indicator Sequence Visualization
To intuitively examine how FinSeer embeds indicator sequences, we use LargeVis (Tang et al., 2016) to reduce the embeddings to two dimensions. In Figure 3, we visualize [^3] https://medviz.org/app/ the vector space of nine indicators, including adjusted close price, open price, high price, low price, close price, volume, movement, returns, and WAP. The results show that the green points, representing movement information, are well-clustered, indicating that FinSeer effectively captures meaningful patterns. However, this also suggests that deeper relationships between stock movements and indicators remain to be explored, highlighting the potential for further research in this area.
Figure 3. Visualization of indicator sequence embeddings on ACL18 dataset.
[…]
5. Related Work
5.1. Stock Movement Prediction
Non-LLM Methods. Traditional approaches to stock movement prediction have focused on various aspects of financial data. One prominent category of methods analyzes stock price sequences and their corresponding technical indicators (Qin et al., 2017; Feng et al., 2018) to identify patterns in historical data for predicting future movements. However, due to the complexity of factors influencing stock prices, subsequent methods have incorporated additional contextual information, such as news articles (Ding et al., 2015; Liu et al., 2018) or social media posts (Xu and Cohen, 2018; Wu et al., 2018). Despite efforts, these methods are highly susceptible to noise and struggle to analyze the vast and diverse nature of financial information effectively.
LLM-based Methods. Recent studies have explored using LLMs for financial prediction tasks, either by fine-tuning open-source models or prompting advanced models like GPT-4. However, even state-of-the-art models, including GPT-4, have achieved only random-guessing-level accuracy in stock movement prediction (Xie et al., 2023b, 2024a, a). This highlights the inherent challenges in identifying and analyzing meaningful patterns in a domain as volatile and multifaceted as stock market prediction.
5.2. Time-series forecasting with LLMs
Non-retrieval Methods. To enhance the performance of LLMs in time-series forecasting, existing methods have primarily focused on aligning temporal and textual data, either by transforming time-series into textual formats or by encoding both modalities into a unified vector space. For instance, a study (Jin et al., 2023) reprograms time-series data into textual representations suitable for LLMs, enhancing prediction accuracy via declarative prompts. Similarly, some other studies (Yu et al., 2023; Liu et al., 2024) explore cross-modal alignment: the former applies LLMs to financial forecasting by integrating stock prices with news data, while the latter introduces a cross-modality framework to align time-series with text for improved predictive performance. While these advancements improve alignment, existing methods often struggle to process large-scale time-series data comprehensively due to LLM input limitations. As a result, effectively leveraging extensive historical data remains a challenge. This highlights the need for retrieval-augmented methods, a gap our approach directly addresses.
RAG Methods. With the rapid advancement of RAG techniques in various applications of LLMs, recent research has begun exploring RAG for time-series forecasting. For instance, TimeRAG (Yang et al., 2024) integrates RAG into time-series forecasting by combining Dynamic Time Warping (DTW) with LLMs to improve prediction accuracy. However, relying solely on numeric similarity is insufficient for financial time-series forecasting, as it fails to capture deeper semantic relationships. This underscores the need for a more targeted retrieval framework tailored to the complexities of financial data, a challenge our framework effectively addresses.
6. Conclusion
In this work, we introduce the first RAG framework for financial time-series forecasting, featuring a novel retrieval mechanism with three key innovations: a fine-tuned 1B parameter LLM (StockLLM), a candidate selection method leveraging LLM feedback, and a training objective that maximizes similarity between queries and historically significant sequences. We also construct new datasets integrating financial indicators and historical stock prices, selected through expert consultation and mutual information analysis, to train the retriever and ensure robust evaluation. Our retriever, FinSeer, leverages these datasets and LLM feedback to uncover deeper patterns beyond numeric similarities. Experiments demonstrate that our RAG framework outperforms bare StockLLM and random retrieval, while FinSeer surpasses existing retrieval methods, achieving an 8% higher accuracy on BIGDATA22 and retrieving more impactful sequences. This work highlights the importance of tailored retrieval models in financial forecasting and provides a robust framework for future research.
Appendix A Data examples
A.1. Examples in the Datastore
An example in our datastore is shown below.
"stock_name":"ABBV",
"query_date":"2014-06-05",
"query_movement":"rise",
"open":54.549999,
"high":55.32,
"low":54.360001,
"close":55.299999,
"adj_close":36.980961,
"volume":4847300,
"MACD_Histogram":0.0621141602,
"macd_crossover":null,
"bollinger_bands":null,
"exceeding_upper":null,
"exceeding_lower":null,
"overbought_and_oversold_conditions":null,
"kdj_crossover":null,
"Returns":0.0131913094,
"VWAP":54.9933333333,
"alpha_smr":0.000549858,
"alpha_mom":0.0121661224,
"alpha002":null,
"alpha006":null,
"alpha009":-0.481476,
"alpha012":-0.481476,
"alpha021":1,
"alpha023":0.0,
"alpha024":-0.769035,
"alpha028":null,
"alpha032":null,
"alpha041":-0.1554335256,
"alpha046":-0.481476,
"alpha049":-0.481476,
"alpha051":-0.481476,
"alpha053":null,
"alpha054":126.4264005641,
"alpha101":-18.282056485
Appendix B Addtional Experimental Results
B.1. Score Calculation of Financial Indicators
We compute the MI scores for each indicator using nonparametric estimation methods provided by mutual_info_regression from the Scikit-learn library. Then we normalize them into range $(0,1)$. These scores provide insight into the strength of the dependency between each financial indicator and the forward return. Higher MI scores indicate stronger relationships, highlighting which indicators are most predictive of future price movements. We select the top-20 indicator groups with the highest MI scores in our candidates, as is shown in Figure 4.
Figure 4. Mutual Information(MI) scores between indicators and forward return.
B.2. Candidate Movement Correlation
For each query with retrieved candidates, we convert the movements of retrieved sequences and LLM’s prediction to numerical representations.
${M_{c_{j}},M_{q}}&=\begin{cases}\text{1},&\text{rise}\\ \text{0},&\text{freeze}\\ \text{-1},&\text{fall}\end{cases}\end{aligned}$
Then, we calculate the average value of the $M_{c_{i}}$ from the five retrieved sequences to represent the RAG-provided result, and the final prediction of the query is recorded as another numerical value:
[…]
Lastly, we adopt the Pearson correlation coefficient to compute the correlation:
[…]